OpenStudy (anonymous):

The rms speed of CH4 is 3.8 * 10^4 cm/s the total kinetic energy of a mixture of 0.8 moles of CH4 and 0.2 moles of C4H10 is?

OpenStudy (anonymous):

@ujjwal

OpenStudy (ujjwal):

masses of CH4 and C4H10 are in the ratio 8:29 so, their rms velocity of C4H10=7.2 * 10^4 cm/s=7.23*10^2 m/s(approx) total KE of mixture=\[\frac{1}{2}[m_{CH4}\times v_{CH4}^2*0.8N_A+m_{C2H10} \times v_{C4H10}^2 \times 0.2 N_A]\] where \(N_A \)=6.022*10^23

OpenStudy (ujjwal):

*the

OpenStudy (anonymous):

@ujjwal the answert should be 1156 J

OpenStudy (ujjwal):

It has lots of weird numbers with decimal.. And i don't feel like calculating the whole thing. Did you solve it? Don't you get the answer?

OpenStudy (anonymous):

let me do it

OpenStudy (anonymous):

i gussed

OpenStudy (anonymous):

i gussed becoz there are many 10^x numbers so how will the answer be correct

OpenStudy (ujjwal):

Do calculations in SI units.

OpenStudy (anonymous):

i got 1676

OpenStudy (anonymous):

*10^24

OpenStudy (anonymous):

wait @ujjwal um how did u get velocity of c4h10

OpenStudy (ujjwal):

That's such a large value.. Either i am wrong or your calculation is wrong.

OpenStudy (anonymous):

ok u used inverse root thing?

OpenStudy (ujjwal):

@A.Avinash_Goutham By using the relation \[v_{rms}=\frac{K}{\sqrt M}\] Ypu already have \(v_{rms}\) of CH4. so, knowing the ratio of masses would do.

OpenStudy (anonymous):

fine @Yahoo! try using this 3/2 nrt and u l get temp by using velocity

OpenStudy (anonymous):

wat is 3/2 nrt

OpenStudy (anonymous):

it's energy of a gas.....

OpenStudy (anonymous):

did nt understand

OpenStudy (anonymous):

um like .5 m v*v = 3/2 nRT

OpenStudy (anonymous):

ok......